OpenStudy (anonymous):

lim (x - sqrt(x^2 + 7)) as x goes to negative infinity. Compute the limit. I just get 0/0 when I go through the process of dividing by the highest power of x in the denominator and accounted for the negative in negative infinity. Is that correct?

OpenStudy (anonymous):

multiply both top and bottom by (x + sqrt(x^2+7))

OpenStudy (anonymous):

Would you like to see the work that i've done? I've already multiplied by the conjugate.

OpenStudy (anonymous):

you should have had -7/(x + sqrt(x^2 + 7) yes?

OpenStudy (anonymous):

yes, the next step being dividing the top and bottom terms by "x" correct?

OpenStudy (anonymous):

actually no, you would not accomplish anything doing so

OpenStudy (jdoe0001):

\(\bf x-\sqrt{x^2}\implies x-x\implies 0\\ \quad \\ x^2+7>x^2\qquad thus\qquad x-\sqrt{x^2+7}\implies \textit{a negative value}\) the bigger "x" becomes, the bigger that negative value I'd think

OpenStudy (anonymous):

wait, nvm

OpenStudy (anonymous):

http://imgur.com/tq6vc3n this is the work i've done after multiplying by the conjugate. i've followed all the rules for infinite limits I believe. Not sure why I'm still ending up getting an indeterminate form. Perhaps an error in the question?

OpenStudy (anonymous):

the problem is that sqrt(x^2) isn't x. This is only true when x > 0

OpenStudy (turingtest):

lim (x - sqrt(x^2 + 7)) as x -> -infty what is the limit as \(x\to-\infty\) of x ? what is the limit as \(x\to-\infty\) of sqrt(x^2+7) ?

OpenStudy (anonymous):

Negative infinity and infinity respectively.

OpenStudy (turingtest):

nice, so that gives the limit as\[-\infty-(\infty)=-\infty\]there is no aspect of the function that does not tend to -infty, hence that's the limit

OpenStudy (anonymous):

Would it still be correct in saying that the limit does not exist since infinity is well, you know, infinite?

OpenStudy (turingtest):

yes, when the limit is infinity you can always say it doesn't exist, because infinity is not a number anyway

OpenStudy (anonymous):

Thank you :]

OpenStudy (turingtest):

welcome!

OpenStudy (anonymous):

the limit is -infinity but how would you show that?

OpenStudy (turingtest):

find the limit of x, then find the limit of -sqrt(x^2+7), then sum the two

OpenStudy (turingtest):

if you got \(-\infty+\infty\), then the limit could be anything, but since *both* parts tend to \(-\infty\), so does the whole thing

OpenStudy (anonymous):

but the whole thing is under the denominator, wouldn't it be 0? -7/(a large number) is closed to 0. But the thing is the limit as x->-inf of 7/(x + sqrt(x^2+7)) is -inf

OpenStudy (turingtest):

I didn't do that conjugate thing, so mine is just as it was originally stated, everything in the numerator

OpenStudy (turingtest):

if you do the conjugate thing it makes it all more complicated, because the you *do* get \[-\infty+\infty\]in the denominator, and that can take on literally *any* value. So in this case, multiplying by the conjugate actually gets us farther from the answer, not closer...

OpenStudy (anonymous):

ahh.. I see